# On the connectedness of the singular set of holomorphic foliations

### v3

@@ Abstract.
Let $\mathscr{F}$ be a singular holomorphic foliation of dimension $k>1$ in a projective $n$-manifold $X$. Assume that the normal sheaf of $\mathscr{F}$ is ample (as is always the case when $X=\mathbb{P}^{n}$) and the singular set $\mbox{Sing}(\mathscr{F})$ has dimension $\leq k-1$. We show that the union of the irreducible components in $\mbox{Sing}(\mathscr{F})$ of dimension exactly k1k-1 is necessarily connected. Consequently, this yields a Bott-type topological obstruction to the integrability of singular holomorphic distributions, echoing Bott’s vanishing theorem, and it answers a question by D. Cerveau for codimension one foliations on $\mathbb{P}^{3}$.
@@ Introduction
In the 1960s, Raoul Bott [3] demonstrated that topological constraints can obstruct the existence of globally integrable foliated structures. In the same vein, we establish a global obstruction to integrability for singular holomorphic distributions. Indeed, there exist holomorphic distributions of dimension $k>1$ on $\mathbb{P}^{n}$, for all $n\geq 3$, whose singular set contains several connected components of dimension k1k-1 (see Remark 4.5). Our main Theorem precludes integrability for all such distributions.
@@
Theorem 1 answers positively the question posed by D. Cerveau in [7] regarding dimension 2 foliations on $\mathbb{P}^{3}$, since the determinant bundle of the normal sheaf of every holomorphic foliation on projective spaces is ample. Cerveau’s problem was motivated by the known models of codimension one foliations classified in degrees 0, 1, and 2.
This falls within the Bott phenomenon of global obstructions to integrability: whereas Bott’s theorem compels the vanishing of invariant polynomials in the Chern classes of the normal bundle in degrees exceeding the codimension for integrable subbundles, our connectedness Theorem 1 furnishes a topological obstruction to integrability for singular distributions under an ampleness hypothesis. We know from [5, Theorem 9.5 item 2b.] and [13, Table 1] that there are families of holomorphic distributions of dimension two on $\mathbb{P}^{3}$, whose singular set contains two connected components of dimension 1.
@@
In [4], the first author, Cerveau, Giraldo, and Lins Neto asked whether the tangent sheaf of a codimension one foliation must split as a sum of line bundles whenever it is locally free. Recently, the third author, Faenzi and Vallés in [12] provided a negative answer to this problem by producing examples of rational foliations on $\mathbb{P}^{3}$ of even degree $2k+4,~k\geq 0$, whose tangent sheaf is a twisted null-correlation bundle, hence non-split; the singular sets of such foliations are singular, non-reduced curves. Furthermore, these foliations admit flat deformations to non-integrable distributions.
@@
Answering the preceding question in the setting of projective threefolds, we show that a natural topological constraint on the tangent sheaf—together with the ampleness of the normal determinant—yields a global obstruction to integrability, detected by the number of isolated singularities of the foliation.
@@
Let $\mathscr{F}$ be a codimension one foliation on a smooth projective threefold $X$ satisfying $h^{1}(\mathcal{O}_{X})=0$, and let $C$ be the union of all 1-dimensional components of ${\rm Sing}(\mathscr{F})$; assume that $C$ is reduced and that the determinant of normal bundle of $\mathscr{F}$ is ample. If $h^{1}(TX\otimes\det(N_{\mathscr{F}})^{*})=h^{2}(TX\otimes\det(N_{\mathscr{F}})^{*})=0$, then
$h^{2}(\mathscr{F}\otimes\det(N_{\mathscr{F}})^{*})=c_{3}(\mathscr{F})\cap[X]$
which corresponds to the number of isolated singularities of $\mathscr{F}$ counted with multiplicities. In particular, when $X=\mathbb{P}^{3}$, the conditions are satisfied and the formula holds.
This is also related to another problem posed by Cerveau in [7, Problem 2.2], which concerns the description of foliations whose singular set has pure dimension one, a condition equivalent to the tangent sheaf being locally free (see [5]). The result provides a topological and algebro-geometric obstruction for such foliations.
In particular, a codimension 1 foliation with locally free tangent sheaf on a smooth 3-dimensional projective manifold, under the hypotheses of Corollary 1, must satisfy
$h^{2}(\mathscr{F}\otimes\det(N_{\mathscr{F}})^{*})=0.$
This condition reflects a topological obstruction for a locally free sheaf to be the tangent sheaf of a foliation.
Moreover, codimension 1 foliations on $\mathbb{P}^{3}$ with locally free tangent sheaf force ${\rm Sing}(\mathscr{F})$ to be singular, unless $\mathscr{F}$ is a pencil of planes.
@@
Let $\mathscr{F}$ be a codimension one foliation on $\mathbb{P}^{3}$. If the tangent sheaf of $\mathscr{F}$ is locally free, then either $\mathscr{F}$ is a pencil of planes, or ${\rm Sing}(\mathscr{F})$ is a singular curve.
We emphasize that the hypothesis that $\mathscr{F}$ is locally free is essential: recall that the singular scheme of a generic rational foliation of type $(a,b)$ consists of a smooth complete intersection curve of degree $ab$ plus
@@
distinct points, see [5, Section 10.1] and [11] for the higher-dimensional case. The example in [12, Theorem 7.1], where the foliation is given by a pencil and the tangent sheaf is a twisted null-correlation, has as singular set a singular curve.
@@ Acknowledgments
CA and JS gratefully acknowledge the University of Bari for its hospitality and support through UNIBA Visiting Professor Program. JS is partially supported by UNAM-PAPIIT grant IN101424. MC is partially supported by the Università degli Studi di Bari and by the PRIN 2022MWPMAB- “Interactions between Geometric Structures and Function Theories” and he is a member of INdAM-GNSAGA; he would like to thank Nigel Hitchin for insightful discussions during his visit to the University of Oxford. MJ is supported by the CNPQ grant number 305601/2022-9, the FAPESP-ANR project number 2021/04065-6.
@@ 1. Proof of Theorem 1
where the last equality follows from equation (1.2). The ampleness says us that here there is an embedding $i:X\hookrightarrow\mathbb{P}^{N},$ for some $N$, such that $\det(N_{\mathscr{F}})=\mathcal{O}_{X}(r)=i^{*}\bigl(\mathcal{O}_{\mathbb{P}^{N}}(r)\bigr).$ Therefore
$c_{1}({\rm det}(N_{\mathscr{F}}))^{n-k+1}=r^{n-k+1}h^{n-k+1}\in H^{2(n-k+1)}(\mathbb{P}^{n})\,$,
where $h:=c_{1}({\mathcal{O}}_{X}(1))$. Hence
$j^{*}_{U_{i}}\big(c_{1}({\rm det}(N_{\mathscr{F}}))^{n-k+1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n-k+1})\in H^{2(n-k+1)}(U_{i})\,.$
Since $k>1$, we can apply $j^{*}_{U_{i}}(h^{k-1})$ in (1.3) and we get
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\!\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n-k+1})\smile j^{*}_{U_{i}}(h^{k-1})\,$,
which implies
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n})\,.$
For each $\alpha$ we have
$\widehat{W_{i\alpha}}]\smile h^{k-1}=P^{-1}[W_{i\alpha}]\smile P^{-1}[W_{i}]=P^{-1}\big([W_{i\alpha}\cap W_{i}]\big)=d_{i\alpha}\,P^{-1}[p_{i\alpha}],$
where $\{p_{i\alpha}\}=W_{i\alpha}\cap W_{i}$ is the (scheme-theoretic) intersection point of $W_{i\alpha}$ with a generic complete intersection
$W_{i}=H_{1}^{i}\cap\cdots\cap H_{n-k+1}^{i},\qquad H_{1}^{i},\dots,H_{n-k+1}^{i}\in H^{0}(X,\mathcal{O}_{X}(1)),$
and
$d_{i\alpha}=\operatorname{length}(W_{i\alpha}\cap W_{i})=(h^{k-1}\!\cdot W_{i\alpha})>0.$
Moreover, $p_{i\alpha}\in U_{i}$ can be defined by hyperplane sections through $p_{i\alpha}$:
$P^{-1}[p_{i\alpha}]=[\,\widetilde{H}^{(i,\alpha)}_{1}\cap\cdots\cap\widetilde{H}^{(i,\alpha)}_{n}\,]=h^{n},\qquad\widetilde{H}^{(i,\alpha)}_{\ell}\in H^{0}(X,\mathcal{O}_{X}(1))\ \text{generic}.$
Thus
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big(d_{i\alpha}P^{-1}[p_{i\alpha}]\big)=r^{n-k+1}\,j^{*}_{U_{i}}(h^{n}),$
and $j^{*}_{U_{i}}(h^{n})=j^{*}_{U_{i}}\big(P^{-1}[p_{i\alpha}]\big)\neq 0$ because $p_{i\alpha}\in Z_{i}\subset U_{i}$. Hence,
(1.4) $\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,d_{i\alpha}=r^{\,n-k+1}.$
Baum–Bott in cohomology gives
$c_{1}(N_{\mathscr{F}})^{\,n-k+1}=\sum_{i=1}^{s}\widehat{\mathrm{Res}}(\mathscr{F},Z_{i})=\sum_{i=1}^{s}\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,[\widehat{W_{i\alpha}}$].
Cupping with $h^{k-1}$ and using $\widehat{W_{i\alpha}}]\smile h^{k-1}=d_{i\alpha}\,h^{n}$ and (1.4), we obtain
$r^{\,n-k+1}h^{n}=c_{1}(N_{\mathscr{F}})^{\,n-k+1}\smile h^{k-1}=\sum_{i=1}^{s}\left(\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}d_{i\alpha}\,h^{n}\right)=\sum_{i=1}^{s}r^{\,n-k+1}h^{n}.$
Hence $h^{n}=s\,h^{n}$, and since $h$ is ample, $h^{n}\neq 0$; therefore $s=1$ and $Z$ is connected.
2. Proof Corollary 1
@@
which is the number of isolated singularities of $\mathscr{F}$ counted with multiplicities, since by [5, Lemma 2.1], the foliation $\mathscr{F}$ has tangent sheaf locally free along $C$. Note that since $C$ is reduced, then it has no embedded points.
@@
$h^{1}(\mathcal{O}_{\mathbb{P}^{3}})=h^{1}(T\mathbb{P}^{2}(-2-d))\simeq h^{1}(\Omega_{\mathbb{P}^{2}}^{1}(2-d))=0\quad\text{for all }d\geq 0.$
@@ 4. Examples and remarks
In this section, we will present examples and observations that demonstrate the necessity of the hypotheses, in particular that Theorem 1 is sharp.
@@
where $\pi_{2}^{-1}(p)=Y$ for all $p\in\operatorname{Sing}(\mathscr{F})$. Then $c_{1}(\det(N(\pi_{2}^{*}\mathscr{F})))=\pi_{2}^{*}c_{1}(\det(N_{\mathscr{F}}))$, and
@@
The next example generalizes to high dimensions an example from [4] for $n=3$; the foliation $\widetilde{\mathscr{F}}_{3,d}$ described below defines a rigid irreducible component of the space of foliations of codimension one on $\mathbb{P}^{3}$ of degree $d+1$, and $\widetilde{\mathscr{F}}_{3,1}$ generates the exceptional component of foliations of degree 2 [4, Cor. 3, p 1007].
@@
Let $\mathscr{F}$ be a holomorphic foliation on $\mathbb{P}^{n}$ of dimension $k>1$, with $Z\subset{\rm Sing}(\mathscr{F})$ the non-empty component of dimension k1k-1. We can write
@@
where $Z_{j}$ are the connected components. In this case, we have
$Z_{i}\cap Z_{k}]=(d_{j}h^{n-k+1})\smile(d_{k}h^{n-k+1})=d_{j}d_{k}h^{2(n-k+1)}\in H^{2(n-k+1)}(\mathbb{P}^{n}),$
with $d_{j}=\deg(Z_{j})$ and $d_{k}=\deg(Z_{k})$. Thus, $Z_{i}\cap Z_{k}]\neq 0$ whenever $n\leq 2k-2$, implying that $Z$ is connected in this case. Consequently, the statement of Theorem 1 is not immediate in the case where n>2k2n>2k-2. Recently, in [10], it was shown that $\mathrm{dim}({\rm Sing}(\mathscr{F}))\geq k-1$, whenever $n\geq 2k$, meaning that $Z_{-}$ is empty.

### v4

@@ Abstract.
Let $\mathscr{F}$ be a singular holomorphic foliation of dimension $k>1$ on a projective $n$-manifold $X$. Assume that the determinant line bundle $\det(N_{\mathscr{F}})$ of the normal sheaf is ample and that the singular set ${\rm Sing}(\mathscr{F})$ has dimension $\leq k-1$. We show that the union of the irreducible components of ${\rm Sing}(\mathscr{F})$ of dimension exactly k1k-1 is necessarily connected. This yields a topological obstruction to the integrability of singular holomorphic distributions, echoing Bott’s vanishing theorem, and it answers a question by D. Cerveau for codimension one foliations on $\mathbb{P}^{3}$.
@@ Introduction
In the 1960s, Raoul Bott [3] demonstrated that topological constraints can obstruct the existence of globally integrable foliated structures. In the same vein, we establish a global obstruction to integrability for singular holomorphic distributions. Indeed, by [5, 13], there exist holomorphic distributions of dimension $k>1$ on $\mathbb{P}^{n}$, for all $n\geq 3$, whose singular set contains several connected components of dimension k1k-1. Our main Theorem precludes integrability for all such distributions.
@@
Theorem 1 answers positively the question posed by D. Cerveau in [7] regarding dimension 2 foliations on $\mathbb{P}^{3}$, since the determinant bundle of the normal sheaf of every holomorphic foliation on projective spaces is ample.
This falls within the Bott phenomenon of global obstructions to integrability: whereas Bott’s theorem compels the vanishing of invariant polynomials in the Chern classes of the normal bundle in degrees exceeding the codimension for integrable subbundles, our connectedness Theorem 1 furnishes a topological obstruction to integrability for singular distributions under an ampleness hypothesis.
@@
In [4], the first author, Cerveau, Giraldo, and Lins Neto asked whether the tangent sheaf of a codimension one foliation must split as a sum of line bundles whenever it is locally free. Recently, the third author, Faenzi and Vallés in [12] provided a negative answer to this problem by producing examples of rational foliations on $\mathbb{P}^{3}$ of even degree $2k+4,~k\geq 0$, whose tangent sheaf is a twisted null-correlation bundle, hence non-split; the singular sets of such foliations are singular non-reduced curves. Furthermore, these foliations admit flat deformations to non-integrable distributions.
@@
We give a partial answer by showing that a natural topological constraint on the tangent sheaf, together with the ampleness of the normal determinant, yields a global obstruction to integrability, detected by the number of isolated singularities of the foliation.
@@
Let $\mathscr{F}$ be a codimension one foliation on a smooth projective threefold $X$ with $h^{1}(\mathcal{O}_{X})=0$, and let $C$ be the union of the one–dimensional components of ${\rm Sing}(\mathscr{F})$. Assume that $C$ is reduced and that $\det(N_{\mathscr{F}})$ is ample. If
$h^{1}(TX\otimes\det(N_{\mathscr{F}})^{*})=h^{2}(TX\otimes\det(N_{\mathscr{F}})^{*})=0,$
then
the number of isolated singularities of $\mathscr{F}$ counted with multiplicities. In particular, when $X=\mathbb{P}^{3}$, these conditions are satisfied, and the formula holds.
This is also related to a problem posed by Cerveau in [7, Problem 2.2], concerning the description of foliations whose singular set has pure dimension one, equivalently, whose tangent sheaf is locally free (see [5]). Our result provides a topological and algebro-geometric obstruction to the existence of such foliations.
In particular, under the hypotheses of Corollary 1, a codimension one foliation on a smooth projective threefold with locally free tangent sheaf must satisfy
$h^{2}(\mathscr{F}\otimes\det(N_{\mathscr{F}})^{*})=0,$
a vanishing that obstructs a locally free sheaf from being the tangent sheaf of a foliation.
Moreover, a codimension one foliation on $\mathbb{P}^{3}$ with locally free tangent sheaf forces ${\rm Sing}(\mathscr{F})$ to be singular, unless $\mathscr{F}$ is a pencil of planes.
@@
Let $\mathscr{F}$ be a codimension one foliation on $\mathbb{P}^{3}$. If the tangent sheaf of $\mathscr{F}$ is locally free, then either $\mathscr{F}$ is a pencil of planes or ${\rm Sing}(\mathscr{F})$ is a singular curve.
The hypothesis that $\mathscr{F}$ has locally free tangent sheaf is essential. Indeed, the singular scheme of a generic rational foliation of type $(a,b)$ consists of a smooth complete intersection curve of degree $ab$ together with
@@
distinct points; see [5, Section 10.1] and [11] for the higher-dimensional case. By contrast, in the example of [12, Theorem 7.1], where the foliation is given by a pencil and the tangent sheaf is a twisted null-correlation bundle, the singular set is a singular curve.
@@ Acknowledgments
CA and JS gratefully acknowledge the University of Bari for its hospitality and support through UNIBA Visiting Professor Program. JS is partially supported by UNAM-PAPIIT grant IN101424. MC is partially supported by the Università degli Studi di Bari and by the PRIN 2022MWPMAB- “Interactions between Geometric Structures and Function Theories” and he is a member of INdAM-GNSAGA; MJ is supported by the CNPQ grant number 305601/2022-9, the FAPESP-ANR project number 2021/04065-6, and the FAPESP-CEPID project number 2024/00923-6.
We extend our deepest thanks to Nigel J. Hitchin on the occasion of his 80th birthday. His mentorship over the years has been a cornerstone of our mathematical lives, and we are forever grateful for the guidance and inspiration he provided as he showed us the beauty found at the intersection of geometry and topology.
@@ 1. Proof of Theorem 1
where the last equality follows from equation (1.2).
Set $L:=\det(N_{\mathscr{F}})$ and write $a:=c_{1}(L)\in H^{2}(X;{\mathbb{Q}})$. Choose $m\geq 1$ such that $L^{\otimes m}$ is very ample and set
$h:=c_{1}(L^{\otimes m})\in H^{2}(X;\mathbb{Z})\subset H^{2}(X;{\mathbb{Q}}).$
Then $a=(1/m)\cdot h$ in $H^{2}(X;{\mathbb{Q}})$. Since $k>1$, cup (1.3) with $h^{k-1}$, viewed in $H^{2(k-1)}(X;{\mathbb{Q}})$ and then pulled back to $H_{c}^{2(k-1)}(U_{i};{\mathbb{Q}})$). We obtain the following equality in $H_{c}^{2n}(U_{i};{\mathbb{Q}})$:
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\!\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=j^{*}_{U_{i}}\!\big(c_{1}(L)^{n-k+1}\smile h^{k-1}\big).$
For each $\alpha$, define
$d_{i\alpha}:=\int_{X}[W_{i\alpha}]\smile h^{k-1}\in\mathbb{Z}_{>0}.$
Equivalently, if $H_{1}^{i},\dots,H_{k-1}^{i}\in|L^{\otimes m}|$ are sufficiently general and
$W_{i}:=H_{1}^{i}\cap\cdots\cap H_{k-1}^{i},$
then $W_{i\alpha}\cap W_{i}$ is a 0–dimensional scheme and $d_{i\alpha}=\mathrm{length}(W_{i\alpha}\cap W_{i}).$ Thus
[…]
where $\eta$ is a fixed generator of the one-dimensional space $H^{2n}(X;{\mathbb{Q}})$. Observe that we do not identify $\eta$ with $h^{n}$; we only use $\eta\neq 0$. Therefore, after evaluating in top degree, we get
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,d_{i\alpha}=\int_{X}c_{1}(L)^{n-k+1}\smile h^{k-1}=\frac{1}{m^{n-k+1}}\int_{X}h^{n}\neq 0.$
In particular, the same nonzero number appears for every $i$. On the other hand, Baum–Bott gives
$c_{1}(L)^{n-k+1}=\sum_{i=1}^{s}\widehat{\mathrm{Res}}(\mathscr{F},Z_{i})=\sum_{i=1}^{s}\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,[\widehat{W_{i\alpha}}]\ \text{in }H^{2(n-k+1)}(X;{\mathbb{Q}}).$
Cupping with $h^{k-1}$ and integrating over $X$ yields
$\int_{X}c_{1}(L)^{n-k+1}\smile h^{k-1}=\sum_{i=1}^{s}\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,d_{i\alpha}=s\cdot\int_{X}c_{1}(L)^{n-k+1}\smile h^{k-1}.$
Since
$\int_{X}c_{1}(L)^{n-k+1}\smile h^{k-1}\neq 0,$
we conclude $s=1$. Hence $Z$ is connected.
2. Proof of Corollary 1
@@
which is the number of isolated singularities of $\mathscr{F}$ counted with multiplicities, since by [5, Lemma 2.1], the foliation $\mathscr{F}$ has tangent sheaf locally free along $C$. Reducedness alone does not exclude embedded associated points in general. In what follows we only use the elementary fact that for a reduced curve $C$, $h^{0}(\mathcal{O}_{C})$ equals the number of connected components of $C$. In particular, if $C$ is reduced and connected, then $h^{0}(\mathcal{O}_{C})=1$.
@@
$h^{1}(\mathcal{O}_{\mathbb{P}^{3}})=h^{1}(T\mathbb{P}^{3}(-2-d))\simeq h^{1}(\Omega_{\mathbb{P}^{3}}^{1}(2-d))=0\quad\text{for all }d\geq 0.$
@@ 4. Examples and remarks
In this section we give examples illustrating the necessity of the hypotheses and the sharpness of Theorem 1.
@@
where $\pi_{2}^{-1}(p)=Y$ for all $p\in\operatorname{Sing}(\mathscr{F})$. Then
$c_{1}(\det(N(\pi_{2}^{*}\mathscr{F})))=\pi_{2}^{*}c_{1}(\det(N_{\mathscr{F}})),$
and
@@
This example generalizes to higher dimensions the example of [4] for $n=3$. The foliation $\widetilde{\mathscr{F}}_{3,d}$ below defines a rigid irreducible component of codimension one foliations on $\mathbb{P}^{3}$ of degree $d+1$, and $\widetilde{\mathscr{F}}_{3,1}$ generates the exceptional component of degree $2$ foliations [4, Cor. 3, p. 1007].
@@
Let $\mathscr{F}$ be a holomorphic foliation on $\mathbb{P}^{n}$ of dimension $k>1$, and let $Z\subset{\rm Sing}(\mathscr{F})$ denote the (non-empty) union of the irreducible components of ${\rm Sing}(\mathscr{F})$ of pure dimension k1k-1. Write
@@
where the $Z_{j}$ are the connected components of $Z$. Assume that $n\leq 2k-2$. Then any two projective subvarieties of $\mathbb{P}^{n}$ of dimension k1k-1 must meet. Indeed, for each $j$ we have $Z_{j}]=d_{j}\,h^{n-k+1}\in H^{2(n-k+1)}(\mathbb{P}^{n})$, where $h\in H^{2}(\mathbb{P}^{n})$ is the hyperplane class and $d_{j}=\deg(Z_{j})>0$. Hence, for any indices $j\neq\ell$,
$\smile[Z_{\ell}]$ $=(d_{j}h^{n-k+1})\smile(d_{\ell}h^{n-k+1})$
$=d_{j}d_{\ell}\,h^{2(n-k+1)}\in H^{4(n-k+1)}(\mathbb{P}^{n}),$
and this class is nonzero because $2(n-k+1)\leq n$. Then, $Z_{j}\cap Z_{\ell}\neq\emptyset$ for all $j\neq\ell$, and in particular $Z$ is connected. Consequently, the content of Theorem 1 is only nontrivial in the range n>2k2n>2k-2. Finally, we note that in the range $n\geq 2k$, it was proved in [10] that $\mathrm{dim}({\rm Sing}(\mathscr{F}))\geq k-1$, so $Z$ is nonempty in that case.